Determine molar mass of gas if its pressure is to fall to one-half of its value in vertical distance of 1 m at 298 K.

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi vishrant,

 ρg(h1-h2)= P-P/2

ρ=PM/RT

h1-h2=1

R=8.314

g=9.8

T= 298

M=RT/2g=8.314*298/2*9.8

M=126.4